# Low-Frequency Black-Box Backdoor Attack via Evolutionary Algorithm

### 23 Feb 2024 v1

@@ 3.3 Frequency Trigger Optimization via Simulated Annealing
Algorithm 1 Searching the optimal trigger in frequency space
@@
2:The optimal frequency trigger $\delta^{*}$
@@
17:return $\delta^{*}\leftarrow\delta_{opt}$
Algorithm 1 describes the workflow of searching our optimal trigger with SA in frequency space. Particularly, it starts with randomly initializing a trigger in frequency domain. After that, SA iteratively improve the trigger stealthiness and effectiveness guided by the objective function in Eq. 7 until meeting the termination criteria (i.e., the temperature T drop to the termination temperature $T_{\emph{f}}$). Specifically, in each iteration, an offspring $\delta^{{}^{\prime}}$ is generated under the constraint in Eq. 7 and injected to the clean image through frequency domain. Meanwhile, frequency band $\nu$ is also randomly altered. Once the new trigger is better than the previous one in terms of the objective value, as calculated in line 4 (for cold start) and 12 (during the iteration) of Algorithm 1 , the new trigger will be adopted, otherwise the above process will be repeated. Finally, the current outperformed optimal trigger among $\delta$ and $\delta_{opt}$ will survive as the new $\delta_{opt}$ and enter the next round. Upon termination, the last $\delta_{opt}$ is the desired optimal trigger that is used by Eq. 4 to produce the poisoned dataset.
@@ 4.1 Experimental Setup
Datasets and Models. Without loss of generality, we evaluate LFBA on five benchmark tasks including handwritten digit recognition on MNIST dataset [27], object classification on CIFAR-10 dataset [25] and Tiny-ImageNet (T-IMNET) dataset [26], traffic sign recognition on GTSRB dataset [21] and face attribute recognition on CelebA [34] dataset. For CelebA, we follow previous works [40, 36] to select the top three most balanced attributes including Heavy Makeup, Mouth Slightly Open, and Smiling. Then we concatenate them to create an eight-label classification task. We test LFBA on both small and large-scale datasets with a wide range of image sizes, including both grayscale and RGB images, to verify attack performance remains consistent across different type of image datasets. Following the previous works [43, 45, 3, 36, 11], we consider various architectures for the image classifier. Specifically, we employ a CNN model [36, 11] for MNIST, PreAct-ResNet18 [18] for CIFAR-10 and GTSRB, and ResNet18 [18] for T-IMNET and CelebA. The details of computer vision tasks, datasets and models are described in Table 4 (Section 0.A.1).
Evaluation Metrics. We evaluate the attack effectiveness based on attack success rate (ASR), i.e. the ratio of poisoned samples successfully misclassified to the target label, and test accuracy (ACC) on clean data for the functionality-preserving requirement. For human inspection, we use PSNR [22], SSIM [50] and LPIPS [52] to evaluate the spatial invisibility between clean and poisoned data. LPIPS utilizes the deep features of CNNs such as AlexNet to identify perceptual similarity, while SSIM and PSNR are calculated based on the statistical pixel-wise similarity.
Implementations. For the default setting, we train the classifiers by SGD optimizer with learning rate $0.01$ and decayed by a factor of 0.1 after every 50 epochs. We set batch size to 64 and total number of epochs to 200. We set $\epsilon$ to 0.1 for MNIST, 0.5 for CIFAR-10, GTSRB, and 1.5 for T-IMNET and CelebA. Following the approach outlined in [42], we select approximately 18.3% of the frequency spectrum in the top-left region to search for our low-frequency trigger. We choose $n=3$ as the number of manipulated frequency bands. For simplicity, we set the poison ratio to only 5% and target label to 7 for all datasets [^2] Our attack is label-independent, i.e., the attacker can easily transfer LFBA attack to any other desired labels by searching the corresponding optimal triggers.. Unless otherwise specified, we use this default setting for LFBA in the following. The detailed device information is shown in Section 0.A.2.
@@ 4.2 Attack Performance
We compare both spatial and frequency attacks including BadNets [15], SIG [1], ReFool [33], WaNet [36], FTrojan [49] and FIBA [12] as baselines to evaluate the effectiveness. Since other backdoor attacks [6, 9, 55, 10, 54] require full control of training process and knowledge of the victim classifiers, we do not consider above methods as practical baselines.
Attack effectiveness. We first demonstrate that our LFBA attack achieves high ASR ($\geq$ 99%) across 5 datasets and 3 models with slight accuracy degradation ($<$0.55% in average) (see Table 1). This fact proved that our attack outperforms other black-box attacks in most tasks. That is so because we approximate the trigger effectiveness during the optimization process of SA (see Eq. 7) whereas other methods do not take into account attack effectiveness when designing their triggers. Hence, our experimental results sound urgent alarms to the physical world: an adversary can compromise any architecture using publicly available images injected by imperceptible triggers in dual space, without access to the victim classifier.
Natural stealthiness. A dual-space stealthiness penalty is added to the process of searching the optimal LFBA trigger to ensure natural stealthiness of the poisoned images. We compare the state-of-the-art invisible attacks in the spatial and frequency domain. For each dataset, we randomly select 500 sample images from test dataset to evaluate the trigger stealthiness. A higher PSNR or SSIM or a smaller LPIPS value indicates a better stealthy of an given poisoned sample. It is worth noting that LFBA achieves superior natural stealthiness than current frequency backdoor attacks (see Table 2) due to the less number of frequency bands and minimal perturbations. Such minor alterations of LFBA in frequency space could naturally provide invisibility to the potential defender who lacks knowledge of the correspondent clean image.
@@ 4.3 Attack Against Defensive Measures
We evaluate attack robustness of LFBA against the mainstream defenses including Neural Cleanse [48], STRIP [13], Fine-pruning [31]. We also show the imperceptible frequency artifacts of LFBA against frequency artifacts inspection [51]. Moreover, we evaluate our attack under image preprocessing-based defenses as in [23, 49] to show the practical robustness in a comprehensive way.
Neural Cleanse (NC). The key intuition of NC is that a backdoor trigger can cause any input misclassified to target label. It first reverses engineering possible triggers and detect backdoors in the victim model using anomaly index. Anomaly index larger than 2 indicates the model has been poisoned. Our attack is below the threshold and successfully evade defense on all datasets (see Figure 2 (a)). The reason is that NC focuses on identifying small, fixed patches. However, our LFBA designs trigger in frequency space, in which inserting an imperceptible frequency perturbation causes a minimal change in the entire pixel domain. Consequently, trigger spans the entire pixel space, providing considerable natural similarity. As a result, it fails to effectively reverse our frequency trigger.
STRIP assumes that the predictions given by a backdoored model on poisoned samples consistently tend to be the target label and are not easily changed. It detects poisoned samples by analyzing the classification entropy after superimposing some random clean samples on the test samples. We can observe that LFBA achieves almost the same entropy probability distributions as clean samples’ (see Figure 3) and can bypass defense. The overlap area of distributions refers to the difficulty of poisoned sample detection. For example, the distributions of clean and posioned samples on CIFAR-10 are almost indistinguisgable which indicates that STRIP is hard to detect our posioned samples. This is because superimposing random images in spatial space would introduce frequency disparities in the poisoned sample. Therefore, the prediction of the superimposed image will also change significantly, similar to clean scenario.
Fine-pruning. Fine-pruning assumes that those neurons are more likely contributed to a backdoor task. Based on this assumption, it mitigates backdoor effectiveness by pruning dominant neurons with very low activations by using a small clean dataset. We test LFBA against Fine-Pruning by showing ACC and ASR with respect to the ratio of pruned neuron on GTSRB, CIFAR-10 and T-IMNET (see Figure 2 (b)-(d)). Across all datasets, we observe that the ASR is always higher than ACC without any degradation, making backdoor mitigation impossible. This suggests that Fine-pruning is ineffective against LFBA.
@@
Figure 2: (a): The results of LFBA under NC on different datasets; (b)-(d): The attack effectives of LFBA against Fine-pruning.
@@
Network Inspection. We further investigate how LFBA affects the attention of the classifier by Grad-CAM [41] and illustrate why LFBA provides attack robustness under existing defenses. Grad-CAM finds the critical regions of input images that mostly activate model’s prediction. We showcase visual heatmaps of several images on GTSRB, CIFAR-10, T-IMNET and CelebA (see Figure 4). We first observe that LFBA only slightly diverts the model’s attention away from the correct regions on all datasets. We can explain it by the fact that our trigger is inserted in low-frequency components, which contain the semantics of images, making network attention unaltered. The results suggest the attack robustness of LFBA against backdoor defenses from network inspection perspective.
@@
Image Preprocessing-based Defenses. We select image preprocessing methods in [49, 23], including Gaussian filter, Wiener filter, BM3D [8] and JPEG compression [47], that directly denoise or compress input images. We apply these operations to poisoned test images with various hyperparameters before inference. The results of attack robustness against preprocessing operations are shown in Table 3. It demonstrates that all transformations are effective to remove trigger effectiveness in FTrojan which handcraft mid and high frequency components with anomaly perturbations. However, LFBA can circumvent these defenses since denoisng transformations or lossy compression do not typically operate on the lower frequency components [16]. We further showcase poisoned images and their frequency disparities (compared to clean images) under preprocessing-based operations in Figure 9 (see Section 0.A.4) and explain why LFBA is more robust than existing attacks against such image transformations.
Table 3: Attack robustness of various triggers against preprocessing-based defenses. To illustrate the robustness of our low-frequency trigger, we introduce a full-spectrum variant of LFBA for comparison, named LFBA-Full, which searchs the trigger across the entire spectrum with same attack settings.
@@
Frequency Artifacts Inspection. We consider the same frequency artifacts inspection method as in [51]. In Figure 5, we compare the frequency spectrum between clean and poisoned images with various spatial and frequency triggers. We can first observe that current spatial backdoors, such as SIG, have larger disparities than frequency backdoors although they could achieve perceptual similarity. Moreover, state-of-the-art frequency backdoors also introduce anomaly frequency artifacts. It is worth noting that two spikes lie in central and bottom-right regions in FTrojan’s spectrum and a noticeable color shift exists in FIBA’s. However, our FLBA spectrums look similar to clean images on both datasets, which provides the same smooth spectrums as clean samples’. According to [2, 44, 51], our poisoned samples exhibit the same frequency properties as natural images due to dual-space stealthiness constraint. Therefore, frequency inspection is ineffective to detect anomaly artifacts of LFBA.
In conclusion, a wide range of experimental results empirically demonstrate that LFBA can elude or significantly degrade the performance of the state-of-the-art defenses in both spatial and frequency space even when both the model and defense strategy are unknown. Additionally, our frequency trigger is resilient to image processing-based defenses which provide a more comprehensive robustness than existing attacks. It indicates that LFBA plays an important role in the security of machine learning systems.
Figure 5: Visualization of DCT spectrums between clean and posioned samples under various spatial and frequency attacks on two different input-space datasets including CIFAR-10 ($32\times 32$) and CelebA ($64\times 64$). We randomly select 10000 samples from each dataset and showcase the averaged spectrum results.
@@ 4.4 Ablation Study
We here analyze several hyperparameters that are critical for the LFBA performance including poison ratio, frequency stealthiness budget $\epsilon$ and number of manipulated bands $n$.
Poison ratio $\rho$. We tested LFBA’s effectiveness across a range of $\rho$ values from 0.1% to 10%, and verified that LFBA is robust to this hyperparameter. Details can be seen in Section 0.A.3.
@@
Frequency stealthiness constraint $\epsilon$ and number of manipulated band $n$. $\epsilon$ restrains the maximum perturbation of each frequency band while $n$ controls the number of manipulated frequency bands. We visualize the impact of $\epsilon$ and $n$ on the poisoned images (see Figure 6). If $\epsilon$ is set too large, the poisoned image may be easily recognized (i.e., no stealthiness) by human inspection in pixel domain. Moreover, this could introduce distinguishable disparities in the frequency space. On the other hand, if we set $\epsilon$ too small, the trigger will have a low proportion of features in dual space. In this sense, the classifier will encounter difficulty in catching and learning these trigger features, resulting in a drop of attack effectiveness. Figure 7 illustrates that how $\epsilon$ and $n$ influence the attack effectiveness among the tasks. The ASRs of LFBA decline seriously and eventually reach below 20% while we keep decreasing $\epsilon$ to $0.01$, in which evidences can be seen in GTSRB under various $n$, while there is a drastic drop happens from $\epsilon=0.5$ to $\epsilon=0.1$ in CIFAR-10. Increasing $n$ enhances the effectiveness of LFBA. For instance, the ASR increases from $77\%$ to $92.8\%$ under $\epsilon=0.1$ when $n$ increases from $1$ to $4$ in GTSRB. We can observe that a large $\epsilon$ allows single injection in frequency band to achieve a high ASR. For example, selecting $\epsilon=1$ and $n=1$ to perform our attack can achieve nearly $100\%$ ASR. However, such an attack setup may compromise frequency stealthiness. Thus, a balance between dual-space stealthiness and attack effectiveness should be considered before conducting a LFBA attack.

### 06 Mar 2024 v2

@@ 3.3 Frequency Trigger Optimization via Simulated Annealing
Algorithm 1 Optimal frequency trigger search via SA
@@
2:The optimal frequency perturbations and their bands $\delta^{*}$, $\nu^{*}$
@@
17:return $\delta^{*}\leftarrow\delta_{opt},\nu^{*}\leftarrow\nu_{opt}$
Algorithm 1 describes the workflow of searching our optimal trigger with SA in low-frequency region $(h_{LF},w_{LF})$. Particularly, it starts with randomly initializing a trigger in frequency domain. After that, SA iteratively improves the trigger stealthiness and effectiveness guided by the objective function in Eq. 8 until reaching the termination criteria (i.e., the temperature T drops to the termination temperature $T_{\emph{f}}$). Specifically, in each iteration, an offspring $\delta^{{}^{\prime}}$ is generated under the constraint in Eq. 8 and injected to the clean image through frequency domain. Meanwhile, frequency band $\nu$ is also randomly altered. Once the new trigger is better than the previous one in terms of the objective value, as calculated in line 4 (for cold start) and 12 (during the iteration) of Algorithm 1 , the new trigger will be adopted, otherwise the above process will be repeated. Finally, the current outperformed optimal trigger among $\delta$ and $\delta_{opt}$ will survive as the new $\delta_{opt}$ and enter the next round. Upon termination, the last $\delta_{opt}$ is the desired optimal trigger that is used by Eq. 5 to produce the poisoned dataset.
@@ 4.1 Experimental Setup
Datasets and Models. Without loss of generality, we evaluate LFBA on five benchmark tasks including handwritten digit recognition on MNIST [27], object classification on CIFAR-10 [25] and Tiny-ImageNet (T-IMNET) [26], traffic sign recognition on GTSRB [21] and face attribute recognition on CelebA [34]. For CelebA, we follow [40, 36] to select the top three most balanced attributes including Heavy Makeup, Mouth Slightly Open, and Smiling. Then we concatenate them to create an eight-label classification task. We test LFBA on both small and large-scale datasets with a wide range of image sizes, including both grayscale and RGB images, to verify attack performance and also remain consistency across different types of image datasets. Following [43, 45, 3, 36, 11], we consider various architectures for the image classifier. Specifically, we employ a CNN model [36, 11] for MNIST, PreAct-ResNet18 [18] for CIFAR-10 and GTSRB, and ResNet18 [18] for T-IMNET and CelebA. The details of computer vision tasks, datasets and models are described in Table 4 (Section 0.A.1).
Evaluation Metrics. We evaluate attack effectiveness based on attack success rate (ASR), i.e. the ratio of poisoned samples successfully misclassified to the target label, and test accuracy (ACC) on clean data for functionality-preserving requirement. For human inspection, we use PSNR [22], SSIM [50] and LPIPS [52] to evaluate spatial invisibility between clean and poisoned data. LPIPS utilizes deep features of CNNs to identify perceptual similarity, while SSIM and PSNR are calculated based on the statistical pixel-wise similarity.
Implementations. For the default setting, we train the classifiers by SGD optimizer with learning rate $0.01$ and decayed by a factor of 0.1 after every 50 epochs. We set batch size to 64 and total number of epochs to 200. We set $\epsilon$ to 0.1 for MNIST, 0.5 for CIFAR-10, GTSRB, and 1.5 for T-IMNET and CelebA. Following the approach outlined in [42], we select approximately 18.3% of the frequency spectrum in the top-left region to search for our low-frequency trigger. We choose $n=3$ as the number of manipulated frequency bands. For simplicity, we set the poison ratio to only 5% and target label to 7 for all the datasets [^2] Our attack is label-independent, i.e., the attacker can easily transfer LFBA attack to any other desired labels by searching the corresponding optimal triggers.. Unless explicitly stated otherwise, we adopt this default setting for LFBA in the subsequent sections. The implementation of LFBA is based on PyTorch [37] and executed on a workstation with 16-core AMD Ryzen 9 7950X CPU, NVIDIA GeForce RTX 4090 and 64G RAM.
@@ 4.2 Attack Performance
We compare both spatial and frequency attacks including BadNets [15], SIG [1], ReFool [33], WaNet [36], FTrojan [49] and FIBA [12] as baselines to evaluate the effectiveness. Since other backdoor attacks [6, 9, 55, 10, 54] require full control over training process and knowledge of the victim classifiers, we do not consider the above methods as practical baselines.
Attack effectiveness. We first demonstrate that LFBA achieves high ASR ($\geq$ 99%) across 5 datasets and 3 models with slight accuracy degradation ($<$0.55% in average) (see Table 1). The results confirm that our attack outperforms other black-box attacks in most tasks. That is so because we approximate the trigger effectiveness during the optimization process of SA (see Eq. 8) whereas others do not take into account attack effectiveness when designing their triggers. Our experimental findings raise urgent concerns for the physical realm: adversaries can compromise any architecture by injecting publicly available images with imperceptible triggers in dual space, even without access to the victim classifier.
Natural stealthiness. A dual-space stealthiness penalty is added to the process of searching the optimal LFBA trigger so as to ensure natural stealthiness of poisoned images. We compare the state-of-the-art invisible attacks in spatial and frequency domains. For each dataset, we randomly select 500 sample images from test dataset to evaluate trigger stealthiness. A higher PSNR/SSIM or a smaller LPIPS value indicates a better stealthiness of an given poisoned sample. LFBA achieves more natural stealthiness than current frequency backdoor attacks (see Table 2) due to its less number of frequency bands and minimal perturbations. Such minor alterations of LFBA in frequency space can naturally provide invisibility to the potential defender who lacks knowledge of the correspondent clean image.
@@ 4.3 Attack Against Defensive Measures
We evaluate attack robustness of LFBA against the mainstream defenses including Neural Cleanse [48], STRIP [13], Fine-pruning [31] and network inspection. We also show imperceptible frequency artifacts of LFBA against frequency artifacts inspection [51]. Moreover, we evaluate our attack under preprocessing-based defenses as in [23, 49] to comprehensively illustrate the practical robustness.
Neural Cleanse (NC). The key intuition of NC is that a backdoor trigger can cause any input misclassified to target label. It reverses engineering possible triggers and detects backdoors in the victim model using anomaly index. An anomaly index exceeding 2 signifies that the model has been compromised. LFBA remains below the threshold and successfully evades the defense across all datasets (see Figure 2 (a)). We recall that NC focuses on small and fixed patches but LFBA designs trigger in frequency space, wherein it inserts an imperceptible frequency perturbation only causing a minimal change in the entire pixel domain. Consequently, trigger spans the entire pixel space, providing considerable natural similarity.
STRIP. It assumes that the predictions given by a backdoored model on poisoned samples consistently tend to be target label and are not easily changed. It detects poisoned samples by analyzing the classification entropy after superimposing some random clean samples on the test samples. We can observe that LFBA achieves almost the same entropy probability distributions as clean samples (see Figure 3), allowing it to circumvent the defense. The overlap area of distributions refers to the difficulty of poisoned sample detection. For example, the distributions of clean and poisoned samples on CIFAR-10 are almost indistinguishable, indicating that it is hard for STRIP to detect our poisoned samples. This is so because superimposing random images in spatial space may introduce frequency disparities in poisoned samples. Therefore, the predictions of superimposed images will also undergo significant changes, resembling the clean case.
Fine-pruning. It mitigates backdoor effectiveness by pruning dominant neurons with very low activations via a small clean dataset. We test LFBA against Fine-Pruning and demonstrate ACC and ASR with respect to the ratio of pruned neuron on GTSRB, CIFAR-10 and T-IMNET (see Figure 2 (b)-(d)). Across all datasets, we see that the ASR is always higher than ACC without any degradation, making backdoor mitigation impossible. This suggests that Fine-pruning is ineffective against LFBA.
@@
Figure 2: (a): The results of LFBA under NC on different datasets; (b)-(d): The attack effectiveness of LFBA against Fine-pruning.
@@
Network Inspection. We further investigate the impact of LFBA on the attention of the classifier by Grad-CAM [41] and explain why LFBA provides attack robustness under existing defenses. Grad-CAM finds the critical regions of input images that mostly activate model’s prediction. We showcase visual heatmaps of images on GTSRB, CIFAR-10, T-IMNET and CelebA (see Figure 4). We observe that LFBA does not introduce anomaly attention area of networks when compared to correct regions across all datasets. This is because our trigger is inserted in low-frequency components, which contain the semantics of images, making network attention unaltered. The results suggest the attack robustness of LFBA against backdoor defenses from a network inspection perspective.
@@
Image Preprocessing-based Defenses. We select image preprocessing methods in [49, 23], including Gaussian filter, Wiener filter, BM3D [8] and JPEG compression [47], which directly denoise or compress input images. We further apply these operations to poisoned test images with various hyperparameters before inference. The results are shown in Table 3. They demonstrate that all transformations are effective to remove trigger effectiveness in FTrojan, which handcrafts mid and high frequency components with anomaly perturbations. However, LFBA can circumvent these defenses since denoisng transformations and lossy compression do not typically operate on the lower frequency components [16]. We also showcase poisoned images and their frequency disparities (compared to clean images) under preprocessing-based operations in Figure 9 (see Section 0.A.3) and explain why LFBA is more robust than existing attacks against such image transformations.
Table 3: Attack robustness of various triggers against preprocessing-based defenses. To illustrate the robustness of our low-frequency trigger, we introduce a full-spectrum variant of LFBA for comparison, named LFBA-Full, which searches the trigger across the entire spectrum with same attack settings.
@@
Frequency Artifacts Inspection. We consider the same frequency artifacts inspection method as in [51]. In Figure 5, we compare the frequency spectra between clean and poisoned images with various spatial and frequency triggers. We can see that current spatial backdoors, such as SIG, have larger disparities than frequency backdoors although they could achieve perceptual similarity. Moreover, state-of-the-art frequency backdoors also introduce anomaly frequency artifacts. It is worth noting that two spikes lie in central and bottom-right regions in FTrojan’s spectra and a noticeable color shift exists in FIBA’s. However, LFBA spectra closely resemble clean images on both datasets (exhibiting similarly smooth spectra as clean samples). According to [2, 44, 51], our poisoned samples exhibit the same frequency properties as natural images due to dual-space stealthiness. Therefore, frequency inspection is ineffective to detect anomaly artifacts of LFBA.
In conclusion, a wide range of experimental results empirically demonstrate that LFBA can elude or significantly degrade the performance of the state-of-the-art defenses in both spatial and frequency spaces even when both the model and defense strategy are unknown. Besides, our frequency trigger is resilient to image preprocessing-based defenses, which provides more robustness than existing attacks. The results also indicate the pivotal role of LFBA in bolstering the security of machine learning systems.
Figure 5: Visualization of DCT spectra between clean and posioned samples under various spatial and frequency attacks on two different input-space datasets including CIFAR-10 ($32\times 32$) and CelebA ($64\times 64$). We randomly select 10000 samples from each dataset and showcase the averaged spectrum results.
@@ 4.4 Ablation Study
We here analyze several hyperparameters that are critical for the LFBA performance including frequency stealthiness budget $\epsilon$, number of manipulated bands $n$ and poison ratio $\rho$.
@@
Frequency stealthiness constraint $\epsilon$ and number of manipulated band $n$. $\epsilon$ restrains the maximum perturbation of each frequency band while $n$ controls the number of manipulated frequency bands. We visualize the impact of $\epsilon$ and $n$ on the poisoned images (see Figure 6). If $\epsilon$ is set too large, the poisoned image may be easily recognized (i.e., lacking stealthiness) upon human inspection in the pixel domain. Moreover, this could introduce distinguishable disparities in the frequency space. On the other hand, setting $\epsilon$ too small results in the trigger having a low proportion of features in dual spaces. In this sense, the classifier will encounter difficulty in catching and learning these trigger features, yielding a drop of attack effectiveness. Figure 7 illustrates the influences of $\epsilon$ and $n$ on the attack effectiveness among the tasks. The ASRs of LFBA decline significantly and eventually fall below 20% as we continuously decrease $\epsilon$ to $0.01$, in which evidences can be seen in GTSRB under various $n$, while there is a drastic drop occurs from $\epsilon=0.5$ to $\epsilon=0.1$ in CIFAR-10. We notice that increasing $n$ enhances the effectiveness of LFBA. For instance, the ASR increases from $77\%$ to $92.8\%$ under $\epsilon=0.1$ when $n$ increases from $1$ to $4$ in GTSRB. We also observe that a large $\epsilon$ allows single injection in frequency band to achieve a high ASR. For instance, selecting $\epsilon=1$ and $n=1$ to perform our attack can achieve nearly $100\%$ ASR. However, such an attack setup could compromise frequency stealthiness. Thus, it’s crucial to consider a balance between dual-space stealthiness and attack effectiveness before conducting a LFBA attack.
@@
Poison ratio $\rho$. We test LFBA’s effectiveness across a range of $\rho$ values from 0.1% to 10%, and verify that LFBA is robust to this hyperparameter. Details can be seen in Section 0.A.2.